OpenStudy (anonymous):

Solve the equation for y. y -3 = 2(y - 5) - 5y +6

OpenStudy (texaschic101):

y - 3 = 2(y - 5) - 5y + 6 -- distribute through the parenthesis y - 3 = 2y - 10 - 5y + 6 -- combine like terms y - 3 = -3y - 4 -- add 3y to both sides y + 3y - 3 = -4 -- add 3 to both sides y + 3y = -4 + 3 -- combine like terms 4y = -1 -- divide by 4 y = -1/4 check.. y - 3 = 2(y - 5) - 5y + 6 -1/4 - 3 = 2(-1/4 - 5) - 5(-1/4) + 6 -1/4 - 12/4 = 2(-1/4 - 20/4) + 5/4 + 6 -13/4 = 2(-21/4) + 5/4 + 6 -13/4 = - 42/4 + 5/4 + 24/4 -13/4 = -42/4 + 29/4 -13/4 = - 13/4 (correct) solution : y = -1/4